# Weyl symmetry without the traceless condition

### v3

@@ VIII Observational consequences of CGR: The redshift measurements
To illustrate the modification of the frequency redshift achieved by the assumption of CGR theory under the active approach to Weyl transformations, let us consider the Pound-Rebka and Pound-Snider experiments pound-1960 ; pound-1964 ; pound-1965 . A target $T_{h}$ made of some atom is placed at some height $h$ on top of a tower on the Earth’s surface (radial coordinate $r=R_{\oplus}+h$, where $R_{\oplus}$ is the mean radius of the Earth) and an identical target $T_{0}$ (made of the same atom) is placed at the bottom of the tower (radial coordinate $r_{0}=R_{\oplus}$.) Atoms in $T_{h}$ are in some excited state, so these can emit photons to return to the ground state. The emitted photons are then absorbed by atoms in the target $T_{0}$, which is located at the bottom of the tower. [^15] We omit experimental details such as, for instance, that the target at the bottom of the tower must be mounted in a centrifuge with controlled rotation speed, or that the temperature of the emitter had to be controlled to high precision, or that the role of the emitter and of the absorber was continuously exchanged, etc. If we expand (75) $\phi(r)=\phi(r_{0})+\phi^{\prime}(r_{0})\Delta r$, where $\Delta r=r-r_{0}=h$, then we get:
@@
where we have introduced the following notation:
$\alpha\equiv\frac{\phi^{\prime}(r_{0})}{2\phi(r_{0})}\,h.$ (79)
@@
$\alpha=\frac{2m^{4}}{R^{4}_{\oplus}}\left(\frac{h}{R_{\oplus}}\right)=1.66\times 10^{-42},$ (80)
where we used the following values of the parameters: the gravitational constant $G=6.67\times 10^{-11}\,\text{m}^{3}/\text{s}^{2}\text{kg}$, the mass of Earth $M_{\oplus}=5.97\times 10^{24}\,\text{kg}$, mean radius of Earth $R_{\oplus}=6.37\times 10^{6}\,\text{m}$, the speed of light in vacuum $c=2.99\times 10^{8}\,\text{m}/\text{s}$, and the distance between the source and the absorber established in the original experiment by Pound and Rebka pound-1960 : $h=22.55\,\text{m}$. The value of $\alpha\sim 10^{-42}$ in (80) is far beyond the experimental capabilities on Earth. For comparison, the value of the redshift measured in pound-1960 was $z=2.5\times 10^{-15}$.
Another illustrative example can be given by the arbitrary gauge choice $\phi(r)=\sqrt{6}M_{\text{pl}}(1-2m/r)^{-1/4}$. In this case,
$\alpha=\frac{m\,h}{2R^{2}_{\oplus}}\approx 1.23\times 10^{-15},$
@@
IX Discussion and concluding remarks
@@
Here we have demonstrated that the very well-known result within the HEP community that local-scale invariance in flat spacetime requires the vanishing of the trace of the matter stress-energy tensor is not universal in curved spacetime. The requirement $T^{(m)}=0$ holds wherein the masses of timelike fields are constant parameters that are not affected by the Weyl recalings. In the case of perfect fluids, this result means that, under WTs, the energy density of the fluid (the same as that of its pressure) transforms as $\rho\rightarrow\Omega^{-3}\rho$. Therefore, the Lagrangian density of matter fields and/or of perfect fluids is not form-invariant under the conformal transformations. In contrast, if the masses of timelike fields are point-dependent quantities transforming as $m\rightarrow\Omega^{-1}m$, under the WTs, and the energy density of perfect fluids transforms as $\rho\rightarrow\Omega^{-4}\rho$, the Lagrangian density of matter is indeed form-invariant under Weyl transformations. In this case, Weyl symmetry does not require vanishing trace of the matter SET. Actually, when the matter field has a nonvanishing mass, which is transformed under (11) as $m\rightarrow\Omega^{-1}m$, the Lagrangian density of the matter fields is form-invariant under WTs (48), so that the following Ward identity takes place:
@@
Therefore, $\phi$-EOM (36), coincides with the trace of the Einstein-type EOM. That is, KG-type EOM (36) is not an independent equation, so it does not contribute to the dynamics of the CGR theory. Our result that under the WTs (48), vanishing of the trace of the stress-energy tensor of matter is not required for Weyl invariance to be a symmetry of theory (37), may be of importance for a revised perspective of the Weyl anomaly capper-duff-1974 ; duff-1977 ; rusos-1984 ; affleck-1986 ; duff-1994 ; odintsov-2000 ; cai-2010 . Unfortunately, the investigation of this issue goes clearly beyond the aims of the present document.
@@
Interestingly, the above Weyl invariant framework may provide a suitable classical description of the gravitational interactions of matter, which meets Dicke’s principle that the (classical) laws of physics must be invariant under the transformations of units; the name used by Dicke for the Weyl transformations (28). Frequent arguments against the Weyl invariance of physical laws must be adequately identified as arguments against local scale invariance. The latter must not be confused with form-invariance under Weyl transformations:
@@
Acknowledgments The author thanks Felipe F Faria for pointing out several bibliographic references to us and for the useful exchange of ideas and critical comments.
Funding: This work has been supported by FORDECYT-PRONACES-CONACYT under grant CF-MG-2558591.
@@
During the preparation of this work, the author used the WRITEFULL tool on overleaf in order to check spelling and grammar. After using this tool, the author reviewed and edited the content as needed and takes full responsibility for the content of the published article.
Appendix A Lagrangian density of matter in terms of Weyl invariant quantities
Here, for illustration, we demonstrate equality (45) for perfect fluids and also for the Dirac fermion. Let us start with the Lagrangian density of a perfect fluid. In this case, the demonstration of equality (45) is a very simple exercise. Let us define the Weyl invariant energy density of the fluid:
$\mathfrak{r}:=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{-4}\rho.$ (86)
Hence, trivially,
${\cal L}_{\text{fluid}}:=-\sqrt{-g}\,\rho\equiv-\sqrt{-\mathfrak{g}}\,\mathfrak{r},$ (87)
where the left-hand Lagrangian density is written in terms of the fields that suffer the WTs, while the right-hand Lagrangian is written in terms of Weyl invariant quantities.
A.1 The Dirac Lagrangian density
The Lagrangian density of a Dirac fermion $\psi$ ($\bar{\psi}$ is the adjoint spinor of the fermion) in curved space is given by:
[…]
where $\cancel{\cal D}:=\gamma^{\mu}{\cal D}_{\mu}$ and the following covariant derivative has been defined:
${\cal D}_{\mu}\psi=\left(D_{\mu}-\frac{1}{2}\sigma_{ab}e^{b\nu}\nabla_{\mu}e^{a}_{\nu}\right)\psi.$ (89)
In these equations $a,b,c=0,1,2,3$ are the flat spacetime indices, while $\gamma^{a}$ are the Dirac gamma matrices, $e^{a}_{\mu}$ are the tetrad fields such that $g_{\mu\nu}=\eta_{ab}e^{a}_{\mu}e^{b}_{\nu}$ ($\eta_{ab}$ is the Minkowski metric). Besides $\gamma^{\mu}=e^{\mu}_{c}\gamma^{c}$, and
$\sigma_{ab}=\frac{1}{2}\left[\gamma_{a},\gamma_{b}\right]=\frac{1}{4}\left(\gamma_{a}\gamma_{b}-\gamma_{b}\gamma_{a}\right),$
are the generators of the Lorentz group in the spin representation. Above, we have used the standard definition of the $SU(2)\times U(1)$ gauge derivative, $D_{\mu}\psi$ (see, for example, Equation (32) of quiros_prd_2025 ).
Let us define the following Weyl invariant quantities:
$\mathfrak{e}^{a}_{\mu}:=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)e^{a}_{\mu},\;\mathfrak{g}_{\mu\nu}:=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{2}g_{\mu\nu},\;\Psi:=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{-\frac{3}{2}}\psi,\;\mathfrak{m}:=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{-1}m.$ (90)
We can also compute the Levi-Cività (LC) connection (Christoffel symbols) of the Weyl invariant metric $\mathfrak{g}_{\mu\nu}$:
$\mathfrak{C}^{\alpha}_{\mu\nu}=\frac{1}{2}\mathfrak{g}^{\alpha\lambda}\left(\mathfrak{g}_{\mu\lambda,\nu}+\mathfrak{g}_{\nu\lambda,\mu}-\mathfrak{g}_{\mu\nu,\lambda}\right),$ (91)
and the related covariant derivative operator $\mathfrak{V}_{\mu}$. We have
$e^{b\nu}\nabla_{\mu}e^{a}_{\nu}=\mathfrak{e}^{b\nu}\mathfrak{V}_{\mu}\mathfrak{e}^{a}_{\nu}+\left(\mathfrak{e}^{a}_{\mu}\mathfrak{e}^{b\lambda}-\mathfrak{e}^{b}_{\mu}\mathfrak{e}^{a\lambda}\right)\frac{\partial_{\lambda}\phi}{\phi},$
so
$\sigma_{ab}e^{b\nu}\nabla_{\mu}e^{a}_{\nu}=\sigma_{ab}\left(\mathfrak{e}^{b\nu}\mathfrak{V}_{\mu}\mathfrak{e}^{a}_{\nu}+2\mathfrak{e}^{a}_{\mu}\mathfrak{e}^{b\lambda}\frac{\partial_{\lambda}\phi}{\phi}\right).$
Let us next define the Weyl invariant gamma matrix in a curved background;
$\Gamma^{\mu}:=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{-1}\gamma^{\mu}=\mathfrak{e}^{\mu}_{a}\gamma^{a}=\mathfrak{e}^{\mu a}\gamma_{a}.$ (92)
It is straightforward to find
$\cancel{D}\psi=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{\frac{5}{2}}\left[\Gamma^{\mu}\left(D_{\mu}-\frac{1}{2}\sigma_{ab}\mathfrak{e}^{b\nu}\mathfrak{V}_{\mu}\mathfrak{e}^{a}_{\nu}\right)\Psi+\left(\frac{3}{2}\mathfrak{e}^{b\lambda}\gamma_{b}-\sigma_{ab}\gamma^{a}\mathfrak{e}^{b\lambda}\right)\frac{\partial_{\lambda}\phi}{\phi}\Psi\right],$
that is,
$\cancel{D}\psi=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{\frac{5}{2}}\cancel{\mathfrak{D}}\Psi,$ (93)
where $\cancel{\mathfrak{D}}\Psi=\Gamma^{\mu}\left(D_{\mu}-\frac{1}{2}\sigma_{ab}\mathfrak{e}^{b\nu}\mathfrak{V}_{\mu}\mathfrak{e}^{a}_{\nu}\right)\Psi$ is the Weyl invariant gauge covariant slash derivative and we have used the identity $\sigma_{ab}\gamma^{a}=3\gamma_{b}/2$. Hence,
$\sqrt{-g}\,\bar{\psi}\cancel{D}\psi=\sqrt{-\mathfrak{g}}\,\bar{\Psi}\cancel{\mathfrak{D}}\Psi.$
The right-hand side (RHS) of this equality is written in terms of Weyl invariant quantities. This serves as an additional demonstration that the left-hand side (LHS) is form-invariant under WTs. It is not difficult to demonstrate that,
$\sqrt{-g}\,m\bar{\psi}\psi=\sqrt{-\mathfrak{g}}\,\mathfrak{m}\bar{\Psi}\Psi.$
Hence, we have demonstrated that the Lagrangian density of the Dirac spinor fulfills the following equality:
[…]
where the LHS is written in terms of the fields that suffer the Weyl transformations, while the RHS is written in terms of Weyl invariant quantities.

### v4

@@ VIII.2 Pound-Rebka-Snider experiments
To further illustrate the modification of the frequency redshift in terrestrial experiments under the assumption of CGR theory and the active approach to Weyl transformations, let us consider the Pound-Rebka and Pound-Snider experiments pound-1960 ; pound-1964 ; pound-1965 . A target $T_{h}$ made of some atom is placed at some height $h$ on top of a tower on the Earth’s surface (radial coordinate $r=R_{\oplus}+h$, where $R_{\oplus}$ is the mean radius of the Earth) and an identical target $T_{0}$ (made of the same atom) is placed at the bottom of the tower (radial coordinate $r_{0}=R_{\oplus}$.) Atoms in $T_{h}$ are in some excited state, so these can emit photons to return to the ground state. The emitted photons are then absorbed by atoms in the target $T_{0}$, which is located at the bottom of the tower. [^21] We omit experimental details such as, for instance, that the target at the bottom of the tower must be mounted in a centrifuge with controlled rotation speed, or that the temperature of the emitter had to be controlled to high precision, or that the role of the emitter and of the absorber was continuously exchanged, etc. If we expand (74) $\phi(r)=\phi(r_{0})+\phi^{\prime}(r_{0})\Delta r$, where $\Delta r=r-r_{0}=h$, then
@@
where
$\alpha\equiv\frac{\phi^{\prime}(r_{0})}{\phi(r_{0})}\,h\ll 1.$ (88)
@@
$\alpha\approx\frac{2m^{4}}{R^{4}_{\oplus}}\left(\frac{h}{R_{\oplus}}\right)=1.66\times 10^{-42},$ (89)
where we used the following values of the parameters: the gravitational constant $G=6.67\times 10^{-11}\,\text{m}^{3}/\text{s}^{2}\text{kg}$, the mass of Earth $M_{\oplus}=5.97\times 10^{24}\,\text{kg}$, mean radius of Earth $R_{\oplus}=6.37\times 10^{6}\,\text{m}$, the speed of light in vacuum $c=2.99\times 10^{8}\,\text{m}/\text{s}$, and the distance between the source and the absorber established in the original experiment by Pound and Rebka pound-1960 : $h=22.55\,\text{m}$. The value of $\alpha\sim 10^{-42}$ in (89) is far beyond the experimental capabilities on Earth. For comparison, the value of the redshift measured in pound-1960 was $z=2.5\times 10^{-15}$. Another gauge choice that is suitable for removing the spacetime singularity, according to bambi-2017-a , is given by (57): $\phi(r)=\sqrt{6}M_{\text{pl}}/(1+m^{2}/r^{2})$. In this case,
$\alpha\approx 2\left(\frac{m}{R_{\oplus}}\right)^{2}\frac{h}{R_{\oplus}}=3.46\times 10^{-24},$
so that experimental check of this gauge is also beyond the current experimental capabilities.
Another illustrative example can be given by the arbitrary gauge choice $\phi(r)=\sqrt{6}M_{\text{pl}}/(1+m/r)^{1/4}$. In this case,
$\alpha\approx\frac{m\,h}{4R^{2}_{\oplus}}=0.6\times 10^{-15},$
@@
IX Quantum-mechanical removal of spacetime singularities
Let us discuss the possibility that quantum-mechanical considerations within the CGR framework may resolve the spacetime-singularity issue; that is, that quantum theory may remove the spacetime singularities that plague classical gravitational theory penrose-1 ; penrose-2 ; penrose-3 ; hawking-1 ; berger ; bosma . According to the previous discussion in sections V and VII, a global gravitational state $\mathfrak{G}(A|k)=\{\phi_{k}(x),g^{(A|k)}_{\mu\nu}$,...\} can be thought as a fourth-dimensional sheet where the whole history of all fields is drawn: the metric tensor $g^{(A|k)}_{\mu\nu}=g^{(A|k)}_{\mu\nu}(t,\vec{x})$, the scalar field $\phi_{k}=\phi_{k}(t,\vec{x})$, and each matter field coupled to gravity $\chi^{(A)}_{i}=\chi^{(A)}_{i}(t,\vec{x})$ (for simplicity we assume that the matter fields are fixed, that is, these are the same in any gauge). Each sheet is labeled by the pair $(A|k)$, i.e., by the index corresponding to the given GR solution $A=1,2,...,N_{s}$, and by the index corresponding to the gauge choice k=1,2,,Nk=1,2,...,N ($N\rightarrow\infty$). Let us call it a “gauge sheet”. [^22] The gauge sheet is just a vague illustration of the actual meaning of a gauge or what we call here a global gravitational state. A clearer illustration must be consistent with the well-posed initial-value formulation of CGR theory. We assume that such a gauge sheet, for example, the one associated to an arbitrary solution “AA” and a gauge choice $\phi_{k}=\phi_{k}(x)$, is a globally hyperbolic spacetime $({\cal M},g^{(A|k)}_{\mu\nu})_{\phi_{k}}$ admitting foliation by Cauchy hypersurfaces $\sum_{t}$, parametrized by a global time function $t$. We may view this globally hyperbolic spacetime as representing the time development of a Riemannian metric on a fixed three-dimensional manifold wald-book . A given classical gravitational phenomenon is well-described by a single gauge sheet within the infinite collection of $N\to\infty$ gauge sheets, but quantum gravity must take into account all of them, since all of the possible gauges of a given solution contribute to the probability amplitude; [^23] Although the probability amplitude of some solution, in particular (90), is expressible as hawking-2 ; hawking-2 ; hawking-3 : $\varphi=\int{\cal D}\phi\,{\cal D}g\,\exp{iS_{\text{cgr}}[\phi,g]},$ where ${\cal D}\phi$ and ${\cal D}g$, are some measures in the space of all scalar fields $\phi=\phi(x)$ and metrics $g_{\mu\nu}=g_{\mu\nu}(x)$, respectively, and $S_{\text{cgr}}[\phi,g]$ is the classical action of vacuum CGR theory (91), for our qualitative discussion it will be more adequate to work with the sum (90).
$\varphi(A)=\lim_{N\to\infty}\sum_{k=1}^{N}\varphi(A|k),\;\;\varphi(A|k)=\frac{\varphi_{0}}{N}\exp iS_{\text{cgr}}[\phi_{k},g_{(A|k)}],$ (90)
where $\varphi(A)$ is the probability amplitude of the solution “AA”, while $\varphi(A|k)$ is the probability amplitude of the gauge “kk” of solution “AA”, $\varphi_{0}/N$ is some normalization constant, and the sum is over all gauges in the conformal class of the GR solution “AA”: $\mathfrak{C}_{\text{conf}}(A)=\{\mathfrak{G}_{\text{gr}}(A),\mathfrak{G}(A|1),\mathfrak{G}(A|2),...,\mathfrak{G}(A|N-1)\}$. For simplicity of the analysis, we consider the vacuum CGR theory. Hence,
$S_{\text{cgr}}[\phi,g]=\frac{1}{12}\int_{{\cal V}_{4}}d^{4}x\sqrt{-g}\left[\phi^{2}R+6(\partial\phi)^{2}\right]+\frac{1}{6}\int_{\partial{\cal V}_{4}}d^{3}x\sqrt{-h}\phi^{2}K,$ (91)
where ${\cal V}_{4}$ is some integration volume, $h_{\mu\nu}=g_{\mu\nu}\pm n_{\mu}n_{\nu}$ is the metric induced on the boundary of the integration volume $\partial{\cal V}_{4}$, which is orthogonal to the unit vector field $n_{\mu}$, and $K=h^{\mu\nu}K_{\mu\nu}$ ($K_{\mu\nu}=h^{\lambda}_{\;\mu}h^{\sigma}_{\;\nu}\nabla_{\lambda}n_{\sigma}$ is the extrinsic curvature tensor of the boundary). The boundary term is introduced to compensate for the second derivatives of the gravitational action in the variational procedure hawking-2 . This term does not modify the conformal form-invariance of the action (91) (see, for example, appendix A of wands-rev ).
Due to form-invariance of the classical action $S_{\text{cgr}}[\phi,g]$ under Weyl rescalings (45), each probability amplitude $\varphi(A|k)$ in (90) contributes not only a same amplitude $\varphi_{0}/N$, but also a same phase $S_{\text{cgr}}$, so that $\varphi(A|1)=\varphi(A|2)=\varphi(A|*)=...=\varphi(A|N-1)$, where $\mathfrak{G}(A|*)$ can be any of the gauges in the conformal equivalence class $\mathfrak{C}_{\text{conf}}(A)$. Therefore,
$\varphi(A)=\varphi_{0}\exp iS_{\text{cgr}}[\phi_{*},g_{(A|*)}].$ (92)
To explain how the spacetime singularity is removed in quantum-mechanical computations, let us consider a specific solution. Here, for illustration, we consider the GR vacuum Schwarzschild solution (53):
$d\hat{s}^{2}=\hat{g}^{(\cancel{0})}_{\mu\nu}dx^{\mu}dx^{\nu}=-\left(1-\frac{2m}{r}\right)dt^{2}+\frac{dr^{2}}{\left(1-\frac{2m}{r}\right)}+r^{2}d\Omega^{2}.$
Any conformal Schwarzschild metric (54)
$ds^{2}=g^{(\cancel{0}|*)}_{\mu\nu}dx^{\mu}dx^{\nu}=\left[\frac{\sqrt{6}M_{\text{pl}}}{\phi_{*}(r)}\right]^{2}\left[-\left(1-\frac{2m}{r}\right)dt^{2}+\frac{dr^{2}}{\left(1-\frac{2m}{r}\right)}+r^{2}d\Omega^{2}\right],$ (93)
is also a solution of the vacuum CGR-EOM (55). For example, in bambi-2017-a , the authors showed that the gauge choice $\phi_{*}(r)=\sqrt{6}M_{\text{pl}}/\sqrt{1+(L/r)^{4}},$ where $L$ is some length scale, leads to the Schwarzschild conformal metric being complete, so the resulting geometry represents a nonsingular black hole. According to (92), the probability amplitude of the Schwarzschild solution may be given by: $\varphi(\cancel{0})=\varphi_{0}\exp iS_{\text{cgr}}[\phi_{*},g_{(\cancel{0}|*)}],$ where $\phi_{*}$ corresponds to the gauge choice made in bambi-2017-a (see above) and $g^{(\cancel{0}|*)}_{\mu\nu}$ is the conformal Schwarzschild metric (93), which is complete, so there will be no singularity issue. Due to conformal form-invariance of the classical CGR action, leading to the same phase contribution to the probability amplitude from any gauge, we can choose any gauge, say $\mathfrak{G}(A|*)$ $\Rightarrow\,S_{\text{cgr}}[\phi_{*},g_{(\cancel{0}|*)}],$ where the singularity is removed. This is the way in which the spacetime singularities are avoided in quantum-mechanical computations.
A final comment on this issue: Within the path integral formulation of gravity, not only those fields that are solutions of the gravitational EOM, like, for example, $\phi_{*}$ and $g^{(\cancel{0},*)}_{\mu\nu},$ matter. Any field configurations that do not satisfy the gravitational EOM are to be considered as well. Notice, however, that the classical action $S_{\text{cgr}}[\phi,g]$ itself is conformal form-invariant regardless of whether the fields satisfy equations of motion or not. That is, any probability amplitudes contribute the same amplitude and phase. Hence, we are free to choose one of the singularity-free configurations and substitute it in (92), even if the fields do not satisfy any equation of motion.
X Concluding remarks
@@
In this document, we have demonstrated that the very well-known result within the HEP community that local-scale invariance in flat spacetime requires vanishing trace of the matter stress-energy tensor is not universal in curved spacetime. The requirement $T^{(m)}=0$ holds wherein the masses of timelike fields are constant parameters that are not affected by the Weyl recalings. In the case of perfect fluids, this result means that, under WTs, the energy density of the fluid (the same as that of its pressure) transforms as $\rho\rightarrow\Omega^{-3}\rho$. Therefore, the Lagrangian density of matter fields and/or of perfect fluids is not form-invariant under the conformal transformations. In contrast, if the masses of timelike fields are point-dependent quantities transforming as $m\rightarrow\Omega^{-1}m$, under the WTs, and the energy density of perfect fluids transforms as $\rho\rightarrow\Omega^{-4}\rho$, the Lagrangian density of matter is indeed form-invariant under Weyl transformations. In this case, Weyl symmetry does not require vanishing trace of the matter SET. Actually, when the matter field has a nonvanishing mass, which is transformed under (10) as $m\rightarrow\Omega^{-1}m$, the Lagrangian density of the matter fields is form-invariant under WTs (45), so that the following Ward identity takes place:
@@
Therefore, $\phi$-EOM (34), coincides with the trace of the Einstein-type EOM. That is, KG-type EOM (34) is not an independent equation, so it does not contribute to the dynamics of the CGR theory. Our result that under the WTs (45), vanishing of the trace of the stress-energy tensor of matter is not required for Weyl invariance to be a symmetry of theory (24), may be of importance for a revised perspective of the Weyl anomaly capper-duff-1974 ; duff-1977 ; rusos-1984 ; affleck-1986 ; duff-1994 ; odintsov-2000 ; cai-2010 . Unfortunately, the investigation of this issue clearly goes beyond the scope of the present document.
[…]
@@
Interestingly, the above Weyl-invariant framework may provide a suitable classical description of the gravitational interactions of matter, which meets Dicke’s principle that the (classical) laws of physics must be invariant under the transformations of units; the name used by Dicke for the Weyl transformations (25). Frequent arguments against the Weyl invariance of physical laws must be adequately identified as arguments against local scale invariance. The latter must not be confused with form-invariance under Weyl transformations:
@@
Acknowledgments The author thanks Felipe F. Faria for pointing out several bibliographic references and for the useful exchange of ideas and critical comments. FORDECYT-PRONACES-CONACYT has supported this work under grant CF-MG-2558591.
@@
During the preparation of this work, the author used the WRITEFULL tool on Overleaf to check spelling and grammar. After using this tool, the author reviewed and edited the content as needed and takes full responsibility for the content of the published article.
Appendix A Curvature reshift as a solution of the null-geodesic equation
If we write the null geodesic equation (39) in FRW spacetime (64) we have that,
$\frac{d\omega}{d\xi}+a\dot{a}\delta_{ij}k^{i}k^{j}=0,$ (99)
where $\xi$ is the affine parameter along the photon’s path and $k^{\mu}=dx^{\mu}/d\xi$ is the four-wave vector, with components: $k^{0}=dt/d\xi=\omega$ (photon’s cyclic frequency) and $k^{i}=dx^{i}/d\xi$. The following expression is satisfied by the wave vector:
$g_{\mu\nu}k^{\mu}k^{\nu}=0\;\Rightarrow\;-\omega^{2}+a^{2}(t)\delta_{ij}k^{i}k^{j}=0.$
If substitute this equation back into (99) we get that,
$\frac{d\omega}{d\xi}+\frac{\dot{a}}{a}\omega^{2}=0,$
or, if recall that $\omega=2\pi\gamma$ ($\gamma$ is the photon’s frequency) and that $\omega=dt/d\xi$ $\Rightarrow\dot{a}\omega=da/d\xi$, then the null geodesic equation can be written in the following way:
$\frac{d\gamma}{d\xi}+\frac{1}{a}\frac{da}{d\xi}\gamma=0.$ (100)
Straightforward integration of this equation leads to the textbook expression for the cosmological redshift:
$\gamma(t)=\frac{\gamma_{0}}{a(t)},\;\gamma_{0}\equiv e^{C},$ (101)
where $C$ is an integration constant. Since this redshift effect is due exclusively to the propagation of photons in a curved background (expanding FRW spacetime), we call it curvature redshift:
$z_{c}=\frac{a(t_{0})}{a(t)}-1\;\Rightarrow\;a(t)=\frac{a(t_{0})}{1+z_{c}}.$ (102)